OpenStudy (anonymous):

a student borrows $3000 at 15% for 9 months. FInd the total amount due using simple interest

OpenStudy (anonymous):

Simple Interest = Principal*rate* time Principal = $3000 Rate = 0.15 (note 15% = 15/100) time (years) = 9/12 = 3/4 years Apply the Formula: Interest = 3000*0.15*(3/4) = $337.50

OpenStudy (anonymous):

Do you have any questions?